Project

General

Profile

2018-05-01-MEMOMIPO

Joint MICE Experiment Management Office and MICE International Project Office

1st May 2018; 15:30 BST
RAL, R1 1st Floor North, CR22

Present: CW, KL, PG, CR, PS

Notes

  1. Introduction: KL
    • Routine meeting, mainly focused on analysis.
  2. Notes of previous meeting and status of actions: All
    • Notes accepted. No actions recorded in last meeting.
  3. Hall status, plans for equipment repatriation: CW
    • CW will circulate the list of all equipment to be returned. Final comments are requested. The list will be signed-off at the EB on Friday 04May18.
    • A new packing crate for the EMR will be made. Boxes for TOFs still exist and will be reused. CW met earlier today with ABr, GW and JB, RH: now clear what will happen with SS magnets. Remaining issue is that accounts will remain open as before. R.Hale will strip stuff out of teh SSs and pack the equipment in a container for shipping.
      • DC cover: will be organised by CW. Action CW
    • New issue is that copper conductor from RR2 to magnet have significant value, so, plan also to try to revover these. Action CW
    • When ISIS takes over they will assume that all unidentified equipment is scrap. We shall need to make sure appropriate action is taken e.g. w.r.t the FC magnets. Action CW
  4. Software and computing status: DR
    • No report this time as DR is not there.
  5. MLCR and computing decommissioning/plan for continued support for analysis: DR
    • No report this time as DR is not there.
  6. Future proofing the MICE WWW site: DR
    • No report this time as DR is not there.
  7. Paper status and plans: CR
    • Emittance measurement paper: Illness has delayed final draft of MICE note and first draft of paper. Expect both next week.
    • LiH scattering: Believe that the alignment issues have been addressed by revising the fiducial-volume cut. Next issue is the deconvolution on MC was n1ot working well enough. Now believe that the uncertainty is within the systematic uncertainty. This will be revisited at the VC.
    • Amplitude evolution: A list has been made of items that need to be dealt with before it is ready for wise people (Boyd and Neuffer).
    • System-performance paper: Ready now to pull this forward. First set of plots planned for analysis meeting in Warwick.
      • Beam-based alignment: Presently conceived as part of the system performance paper.
    • LH2-scattering paper: Will come when LiH paper is completed. Ggo student has begun to work on this.
    • Field-on scattering: A.Young has begun to work on this. V.Pec has begun on track extrapolation to the absorber.
    • Long-emittance paper: Beginning to work on organising the work on the long paper.
    • PS asked about the publicity for the cooling result. KL had contacted STFC publicity manager and had been told the best approach would be to publicise at the time of the publication. We discussed possible action for Friday and decided not to pursue since time was too short.
  8. AOB
    • Since he could not be at Friday's EB, PS raised:
      • The Common Fund: we agreed that PS would provide a two-pager for submission to the next PSAG to close out the Common Fund. Action PS.
      • UK representation for the EB; we agreed that PS will contact UK group leaders to ensure approprate representation continues. Action PS.

Summary of actions:

  • CW: Organise DC cover for equipment-recovery work in teh MICE Hall.
  • CW: Include copper cables between magnets and RR2 in the scrappage plan.
  • CW: With others make inventry of what may not be scrapped by ISIS post June 2018 when ISIS take over ownershp of the MICE Hall.
  • PS: Generate two-pager to close-out common fund for next PSAG.
  • PS: Close discussion with UK group leaders over UK representation at the MICE EB.

Back to MEMO:

Back

Back to MIPO:

Back

Documents